Grade 11Physical Chemistry

which of the following have minimum wavelength
  1. electron
  2. proton
  3. CO2
  4. SO2

gaurav
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
As we know
\tiny \lambda = h/mv
Now
Which has highest mass will haveeast wavelength.
Now
SO2 = 32 + 32 = 64
CO2 = 12 + 32 = 44
Electron = 9.1 *10^-31
Proton = 1.6 * 10^-27
Hence SO2 will have least wavelength.